16) The hydrostatic pressure at 10 meters depth in the ocean is different than the hydrostatic pressure at 10 meters in a fresh water lake such as Bear Lake. Assuming the ocean is filled with ocean water and that Bear Lake is filled with fresh water, what is the magnitude of the pressure difference at a depth of 10 meters? a) 2.84× 103Pa b) 11.3 × 103 Pa c) 29 Pa d) 14.1×103 Pa
Answer #1

Hydrostatic pressure = P = rho*g*h

where (rho) is the density of the fluid,
g is the acceleration of gravity
h is the height of the fluid above the object

Pressure difference = [rho(ocean water) - rho(fresh water)]*g*h

= 29*9.8*10 = 2.84*10^3 Pa (Option A)